Shift Summary - Evening - 20 june 1999

Robert Tschirhart (tsch@fnal.gov)
Mon, 21 Jun 1999 00:24:00 -0500

Shift Summary for evening shift, 20/jun/99

On Shift: Herman White and Bob Tschirhart

Hours of beam delivered at all intensities: 7.0 hours
Hours of beam delivered at nominal intensity: 7.0 hours
Hours of E832 data (special runs not included): 7.0 hours
Hours of special runs (laser scans, muon runs...): none.

Accounting of lost time: TeV problems during last 1.5 hours of shift.

========================================================================
Runs ended this shift:
Run 13670 - 19.8M events - BEAM_832 - 28.6K vacuum 2pi0

Tapes brought to FCC this shift:, none
(including tapes for online split)

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Took over running experiment.

A few CsI channels show sporadic laser RMS excursions beyond 5%, nothing
permanent.

We tuned up the targeting on upstream SWIC NM2WC1 back to nominal
values. Returning NM2WC1 back to nominal also improved the
position on SEED2.

We ran all the monitoring kumacs, found the following problems:

* 2V TDC shows a long low-side tail, not seen in '97 running.

* DCOR 1Y and 2Y have large inefficiency compared to 1X and 2X.

* CSI channels 449, 455, 2014, and 2274 showed sporadic excursions above
5% in the laser RMS distributions.

* HA system is asking for a HV change to raise gains.

* Two intermittent HCC bits found CsI channels 1534, 2080. Sydney and
Jason notified.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13670 291 18:40 Noticed NM2WC1 targeting off
by 3mm. Stop triggers while
MCR tunes.

22:00 Jim Graham calls, and reminds
us to keep ONSPLIT window active.
Disks 1e, 2e, and 4e getting full.

13670 22:16 Run ended. TeV problems, we
setup for a muon run. Magnet
polarity swapped from pos to neg.

========================================================================~
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Muon run.

Page Elizabeth if any problems in running new regmon kumac during
muon run.